40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 995     Entry time: Fri Sep 26 00:19:54 2008
Author: Jenne 
Type: Update 
Category: PSL 
Subject: Filter-action with the PMC 
Written, but not posted on 24Sept2008:



PMC adventures for this evening



Today's mission was to make more progress on increasing the bandwidth of the PMC servo.



First order of business was to improve the performance of the 14.6kHz notch that Rana put in the PMC servo board a few weeks ago to remove the 14.6kHz body mode resonance of the PMC. Looking at the zoomed in TF that I posted Monday (elog #978), we see that there is still a remnant of a peak near 14.5kHz. A first gut-reaction is that the notch is not tuned properly, that we have just missed the peak. As previously noted in the elog, the peak that we are trying to notch out is at 14.68kHz (elog #874). By unlocking the PMC and measuring the transfer function between FP2 and OutMon (OutMon is the monitor for the high voltage going to the PMC's PZT), I measure the transfer function of the notch, and find that it is notching at 14.63kHz. So we're a teensy bit off, but the Q of the notch is such that we're still getting improvement at the peak frequency. After checking that we are hitting the correct frequency, I put a short (just some wire) around R21, which is the R in the RLC notch filter, to increase the depth of the notch. At the peak frequency of 14.68kHz, we see a 2.5dB improvement of the notch. At the actual notch frequency of 14.63kHz, we see a 3.2dB increase in the depth of the notch. So, shorting R21 helped a little, but not a lot. Also, it's clear that we don't get that much more improvement by being on the resonant frequency, so there's no need to go in and tune the notch on the board.



Second order of business was to investigate the 18.34kHz peak in the transfer function. (Rana spent some time Monday night measuring this peak, and determined that it was at 18.34kHz) We decided that the best plan was to re-implement the Pomona Box notch filter that had previously existed to remove a higher frequency body mode, but tuned for the 18.34kHz mode. I am still not entirely sure what this mode is, but clearly it's a problem by about 20dB (on the TF, the next highest peak is 20dB below the 18.34kHz peak). Unfortunately, while the components should, by Matlab calculations, give me an 18.3kHz notch, I ended up with something like a 21.7kHz notch. This notch is approximately -30dB at 21.7kHz, and -20dB at 18.3kHz. I still need to take transfer functions and power spectra of the PMC servo with this new filter in place to (a) confirm that it did some good, and (b) to determine how important it is that the notch be right-on. More likely than not, I'll take the filter out and fiddle with the capacitors until I get the correct notch frequency.



Third on the list was to lock everything back up (FSS, PMC) after my tinkering, and see what kind of gain we get. Rob and I fiddled with the PMC gain, and it looks like the servo oscillates just before we get up to the max slider gain of 30dB. Looking at the power spectra in DTT, we do not see any significant peaks that suggest oscillation, so it is likely that there is some investigation to be done at frequencies above the 7kHz that we were able to look at with DTT (which isn't surprising, since all of this work has been at 14kHz and higher).



A final note is that we see a feature around 9kHz in the transfer function, and it is not at all clear where it comes from. At this time, it does not seem to be the dominant feature preventing us from increasing the gain, but at some point if we want the bandwidth of the PMC servo to be 10kHz, we'll have to figure this one out.



Still on the PMC todo list:


  • Measure the new transfer function, see if 18.34kHz peak is reduced

  • Tune Pomona Box notch filter to 18.3kHz instead of the current 21.7kHz

  • Retake power spectra of different items on top of PMC, compare to see if there is any one configuration that it obviously better than the others.

  • Find out why the PMC still oscillates when we try to take it up to the max slider gain, and fix it.

PS, is anyone else having trouble getting to the elog from laptops on other parts of the Caltech network (but not LIGO network)? My laptop won't go to the elog, but I can get to the rest of the internet using the Caltech wireless. My computer stopped seeing the elog on Tuesday or so. Joe, do you have any inspiration? Thanks.
ELOG V3.1.3-